OpenStudy (anonymous):

perform the indicated operations (give your answer in scientific notation). (5x10 a small 2 on top by the 10)x(5.6x10 a small 4on top by the 10)

OpenStudy (unklerhaukus):

\[(5\times10^2)\times(5.6\times 10^4)\]\[=(5\times5.6)\times(10^2\times 10^4)\]\[=5\times5.6\times10^{2+4}\]\[=\]\[=\]